# Maritime Cybersecurity: A Comprehensive Review

### v2

@@ 4.1.3. Lesson Learned Summary
Two significant initiatives have emerged to document cybersecurity threats in the maritime sector. Researchers at NHL Stenden University in the Netherlands have compiled the Maritime Cyber Attack Database (MCAD) (O’Dwyer, 2023), a comprehensive archive cataloging over 160 discrete cyber attacks on the maritime transportation sector. Concurrently, Advanced Dataset of Maritime cyber Incidents ReleAsed for Literature (ADMIRAL) has developed a more extensive collection, centralizing data on over 470 cybersecurity incidents within the maritime sector.
@@
Figure 6. Maritime cyber incidents by year 2014 - 2023. Recreated by author based on data from MCAD (https://maritimecybersecurity.nl/) and ADMIRAL dataset (https://www.m-cert.fr/admiral/index.html). Each year’s data is selected from the larger of the two sources.
As shown in Table 2, maritime cyber incidents have affected numerous countries and regions worldwide, including the United States, Europe (e.g., UK, Germany), Asian-Pacific (e.g., Singapore, Japan, Australia), and Middle East (e.g., UAE). This global distribution of attacks underscores that maritime cybersecurity is an international concern, with no country or region immune to these threats. The cyber attacks documented span a range of categories, including abuse and theft of data, DoS attacks, GPS spoofing and jamming, malware infections (particularly ransomware), social engineering and phishing attempts, targeted attacks, and exploitation of system vulnerabilities. This attack range shows how diverse the cyberthreats that the maritime industry faces are. These attacks are carried out by a similarly varied array of players, including nation-states, hacktivists, cybercriminal organizations, and lone hackers. The threat actors’ identities remained a mystery in numerous instances. This diversity of perpetrators suggests that maritime cyberattacks are motivated by a variety of factors.
@@
Table 2. Evaluation of Real-Life Maritime Cyber Attack Incidents
@@
2016 Software Flaws in U.S. Ports USA Shore Website SQL injection attack Ethical Hacker Personal Entertainment View/Modify/Delete Operational Data Integrity
@@
2017 Maersk NotPetya Attack Ukraine Shore Network Malware Advanced Persistent Threat (APT) Groups Political Objectives Financial Loss & Data Contamination Confidentiality
@@
2018 Port of Barcelona Attack Spain Shore Internal IT Systems Ransomware Unknown N/A Cargo Delay Availability
@@
2021 Death Kitty Ransomware South Africa Shore Network Ransomware Cybercriminal Groups Financial Gain Operations Halted & Website Down Availability
2021 Attack on Greek Companies Greece Shore Communication System Ransomware Cybercriminal Groups Financial Gain Data Loss Availability
@@
4.2. Cyber Attack Demonstrations (Academic)
@@
Academic cyber attack demonstrations typically take place in simulated environments or self-developed software that replicate maritime systems and networks. There is a portion of academic research focuses on theoretically feasible attacks by modeling maritime systems to identify potential attack vectors, even if they have not been practically implemented in a full-scale simulation. These demonstrations often focus on specific components of maritime systems (See Table LABEL:tab:summary_threats).
@@
AIS Threats Overview Balduzzi et al. (Balduzzi et al., 2014) conducted a comprehensive security evaluation of the AIS and categorized these threats into three main types: i) spoofing, ii) hijacking, and iii) availability disruption attacks. Kessler et al. (Kessler et al., 2018) presented a comprehensive analysis of the potential cyber threats to the AIS by drawing upon previous research (Gauthier and Seker, 2018)(Purton et al., 2010)(Strohmeier et al., 2014) that examined the vulnerabilities of Automatic Dependent Surveillance-Broadcast (ADS-B), a similar system used in the aviation industry to provide situational awareness for aircraft. Amro and Gkioulos (Amro and Gkioulos, 2022) discussed a novel cyber attack that utilizes the AIS as a covert channel for sending command and control messages and delivering malware to maritime systems.
@@
VDR Threats Overview Harish et al. (Harish et al., 2022) launched a multi-pronged attack on a commercially available VDR, employing a USB Rubber Ducky for physical infiltration, the Metasploit framework for system exploit, a ransomware simulator to demonstrate data encryption risks, and Nmap to identify exploitable network vulnerabilities. Hopcraft et al. (Hopcraft et al., 2023) highlighted cybersecurity deficiencies in current VDR systems and standards, and presented they are susceptible to ransomware attacks, malware infections, denial-of-service (DoS) attacks, insider threats and data breaches. Söner et al. (Söner et al., 2023) used failure modes and effects analysis (FMEA) to assess cyber vulnerabilities and potential attacks against VDR. Data acquisition unit (DAU) is found to be the most vulnerable VDR component, followed by remote access solutions and the bridge control panel. False data injection, command injection, and viruses are ranked as the top three threats in the FMEA risk assessment.
@@
RADAR Threats Overview Junior et al. (Leite Junior et al., 2021) demonstrated how the attacker could send the command to the ship’s radar via an electronic attack (EA) that generates false radar echoes in a specific pattern. Wolsing et al. (Wolsing et al., 2022) presented a taxonomy of possible network cyber attacks on marine RADAR and a simulation environment to implement these attacks. The taxonomy covers denial of service attacks, attacks involving basic image transformations like scaling, rotation and translation, and targeted object manipulations like addition, removal and relocation of radar echoes. Longo et al. (Longo et al., 2023c) demonstrated that it is feasible to develop malware that can autonomously and stealthily target radar systems on ships by exploiting vulnerabilities in standard protocols like NMEA and ASTERIX that lack authentication and encryption. Longo et al. (Longo et al., 2023a) discussed a novel threat model that combines cyber attacks with cyber false flags to target maritime RADAR systems. The proposed cyber attacks manipulate RADAR data to simulate the effects of electronic countermeasures (ECM) like barrage jamming, spot jamming, digital radio frequency memory, and blip enhancement. By making the cyber attacks appear as ECM originating from another ship, it creates a false flag that misleads attribution of the attack. These false flag attacks aim to disrupt radar operations, divert blame, hide the cyber nature of the attack, and project the attacker’s offensive capabilities.
@@
VSAT Threats Overview Santamarta (Santamarta, 2014) discussed security vulnerabilities found in various widely deployed satellite communications (SATCOM) terminals and described several real-world attack scenarios showing how these flaws could be exploited to leak sensitive military data, interfere with distress communications, spoof navigation data, disable safety systems, and more. Based on the two scenarios presented, VSAT systems in the maritime context face threats of data manipulation, spoofing, and communication disruption. Pavur et al. (Pavur et al., 2020) revealed serious security flaws in how maritime VSAT networks are currently implemented. The insecure VSAT links could allow attackers to modify important data like ECDIS navigational charts and AIS vessel tracking information, impacting ship safety. Under certain conditions, an attacker could hijack active TCP sessions and modify data in real-time.
@@
GNSS Threats Overview Grant et al. (Grant et al., 2009) conducted practical experiments to validate the threat of GPS jamming on maritime navigation. Bhatti and Humphreys (Bhatti and Humphreys, 2017) analyzed and demonstrated the ability of an attacker to control a maritime surface vessel by broadcasting counterfeit civil GPS signals. Medina et al. (Medina et al., 2019) carried out a dedicated jamming measurement campaign on the Baltic Sea in cooperation with the German Federal Network Agency.
@@
ECDIS Threats Overview All five papers written by Svilicic et al. (Svilicic et al., 2019a, b, c, 2020, d) focused on the cyber security aspects of the ECDIS and used similar vulnerability scanning methods. However, each paper had a unique focus. Svilicic et al. (Svilicic et al., 2019a) presented an estimation of the cyber security vulnerabilities of ECDIS that arise from weaknesses related to the underlying operating system. Svilicic et al. (Svilicic et al., 2019b) introduced a framework for assessing cyber risks of the ship’s critical systems and assets. Svilicic et al. (Svilicic et al., 2019c) identified the main threats to the ECDIS system through a survey of the ship’s crew and computational vulnerability scanning using Nessus Professional software. Svilicic et al. (Svilicic et al., 2020) highlighted that interconnecting multiple ECDIS workstations with the same vulnerabilities provides an ideal environment for malware to spread. Svilicic et al. (Svilicic et al., 2019d) emphasized on the cyber security weaknesses of a ECDIS arising from the software’s third-party components. Dyryavyy (Dyryavyy, 2014) discussed the cyber security risks and weaknesses within ECDIS and the research revealed that a malicious actor could potentially browse, download, modify, or erase any file on the computer running the ECDIS software (Hayes, 2016).
@@
GMDSS Threats Overview While specific GMDSS cyber attack examples are limited, the concepts of jamming, spoofing and DoS (Akpan et al., 2022)(Tam and Jones, 2018) discussed could all theoretically be used to disrupt GMDSS functionality as ships become more networked and vulnerable to cyber threats. The authors (Tam and Jones, 2018) also discussed the risk of malware infections in ship systems. While not specifically mentioning GMDSS, such an attack could potentially affect GMDSS equipment if it’s connected to infected systems.
Table 3. Summary of the Identified Threats in Academic
Asset Maturity Type Threat Impact Found in Real-Life Maritime Incidents?
AIS Experimental Spoofing Ship spoofing Obscuring illegal activity
[…]
(Balduzzi et al., 2014) Collision spoofing Unnecessary course changes
[…]
Hijacking AIS Hijacking Misleading other vessels
[…]
Theoretical System Failure Device malfunction Incorrect or missing data
Device powered down Vessel disappears
[…]
Radio transmission bit errors Incorrect AIS data
[…]
Network Attacks Man-in-the-middle Eavesdropping, data manipulation and impersonation
[…]
ECDIS Experimental Software Vulnerability Outdated ECDIS software System crashes and unauthorized access
(Dyryavyy, 2014)(Svilicic et al., 2019a) Underlying OS Data manipulation
[…]
Physical Access Unauthorized physical access System tampering, or data theft
Supply Chain Compromised chart updates Navigational errors
[…]
VSAT Experimental Network Attacks Eavesdropping Compromise of personal privacy and business secrets
[…]
RADAR Experimental PPI Poisoning Ghost ship attack Induce wrong decisions
[…]
Jamming Barrage jamming Decrease signal-to-noise ratio
Spot jamming Disrupt radar detection
[…]
GNSS Experimental Spoofing GPS spoofing Incorrect position information
(Bhatti and Humphreys, 2017)(Grant et al., 2009)(Medina et al., 2019) Jamming GPS jamming Loss of GPS positioning
GMDSS Theoretical Jamming Distress alert disruption Prevent distress alerts
[…]
Availability Disruption Denial-of-Service (DoS) Prevent distress alerts
[…]
5. Security Solutions
5.1. Integrated or Holistic
Rather than focusing on individual components, integrated solutions aim to create a comprehensive framework that protects and optimizes the vessel as a whole. For example, an integrated solution might include a centralized security operations center that monitors and manages threats across all vessel systems, from navigation and communication to cargo management and environmental controls. Integrated solutions frequently leverage advanced technologies like machine learning algorithms to analyze data from various shipboard systems, identifying patterns that could indicate potential security breaches or safety risks before they escalate into serious problems.
5.1.1. Machine/Deep Learning-Based Defence
The demonstrated performance and adaptability of some advanced deep learning techniques suggest significant potential for their practical application in safeguarding critical maritime infrastructure against the ever-evolving landscape of cyber threats. In (Kumar et al., 2021), it was revealed that the DLTIF framework exhibited superior performance in automatically extracting and identifying cyber threat patterns compared to traditional machine learning approaches across multiple evaluation metrics. Dual Stack Machine Learning (S2ML) framework is proposed in (Ali et al., 2022) which takes a novel approach by leveraging entropy-based features extracted from network traffic, a method that proves particularly effective in identifying anomalies indicative of DDoS attacks. Similarly, the authors of (Gyamfi et al., 2022) leveraged the capabilities of an adaptive incremental passive-aggressive machine learning (AI-PAML) which is designed to continuously update its learning model as new network attacks are detected.
While DLTIF, S2ML and AI-PAML focus on identifying or detecting threats, the Multi-Agent Reinforcement Learning (MARL) (Wilson et al., 2024) approach takes the next step by exploring how AI can be used to actively defend against these threats. The MARL solution establishes a simulation environment for training autonomous cyber defence agents for maritime operational technology (OT) to enhance the resilience of maritime control systems.
5.1.2. Cryptographic Schemes
The distributed nature of maritime systems, with ships, ports, and other infrastructure being geographically dispersed, renders centralized security solutions impractical. Recent research has proposed several cryptographic schemes that provide security in a distributed manner, employing techniques such as symmetric encryption, attribute-based encryption, erasure coding, identity-based encryption, and digital signatures to enable each entity to safeguard its own data and communications. A lightweight authentication protocol (Chaudhry et al., 2021) using symmetric cryptography primitives like XOR and hash operations can provide mutual authentication and shared key establishment while ensuring vessel privacy. Similarly, an attribute-based secure data aggregation scheme (Wang et al., 2021) employs constant attributes of maritime terminals for authentication and onboard sensors for data encryption, using zero-knowledge proofs for member certification. Addressing a critical aspect of maritime data security overlooked by the previous schemes, it was proposed that a flexible integrity checking and recovery mechanism (Liu et al., 2021) is particularly valuable in maritime environments where harsh conditions frequently lead to data corruption. Other prevention methods include an identity-based authentication mechanism (Gupta et al., 2021) that integrates both cryptographic techniques and distributed ledger technologies to ensure not only the confidentiality and integrity of data but also enhance transparency and non-repudiation in maritime communications.
The prevalence of legacy equipment in maritime communication channels, often predating modern security protocols and persisting due to extended lifecycles, presents significant cybersecurity vulnerabilities in today’s interconnected maritime environment. In order to offer a pragmatic strategy for a gradual, cost-effective transition towards more secure maritime communication paradigms, the authors of (Hemminghaus et al., 2021) proposed the SIGMAR framework, which extends the IEC 61162-450 protocol and integrates the Elliptic Curve Digital Signature Algorithm (ECDSA) to provide robust authentication and integrity verification with minimal latency and communication overhead.
5.1.3. Blockchain-Based Defence
A secure communication framework can be harnessed by introducing a private blockchain network integrated with a terrestrial fusion center for authentication purposes (Rahimi et al., 2020). The approach is designed to provide a resilient and trustworthy communication infrastructure for maritime systems, with particular emphasis on unmanned aerial vehicle (UAV)-assisted sensing applications in maritime environments. In (Zhang et al., 2022), a blockchain-based authentication mechanism is utilized to store and validate the identities of vessels and IoT devices. The solution implements a proof-of-authority consensus mechanism and incorporates a threshold-based approach for detecting malicious nodes. Further, blockchain can be used to secure positioning data in maritime environments. A blockchain-based privacy preserving position data sharing approach is leveraged in (Gai et al., 2022) to address the critical privacy challenges in maritime IoT systems.

### v3

@@ III-A3 Lesson Learned Summary
Two significant initiatives have emerged to document cybersecurity threats in the maritime sector. Researchers at NHL Stenden University in the Netherlands have compiled the Maritime Cyber Attack Database (MCAD) [84], a comprehensive archive cataloging over 160 discrete cyber attacks on the maritime transportation sector. Concurrently, Advanced Dataset of Maritime cyber Incidents ReleAsed for Literature (ADMIRAL) has developed a more extensive collection, centralizing data on over 470 cybersecurity incidents within the maritime sector.
@@
Figure 6: Maritime cyber incidents by year 2014 - 2023. Recreated by author based on data from MCAD (https://maritimecybersecurity.nl/) and ADMIRAL dataset (https://www.m-cert.fr/admiral/index.html). Each year’s data is selected from the larger of the two sources.
@@
TABLE II: Evaluation of Real-Life Maritime Cyber Attack Incidents
@@
2016 Software Flaws in U.S. Ports USA Shore Website SQL injection attack Ethical Hacker Personal Entertain View/Modify Operational Data Integrity
@@
2017 Maersk NotPetya Attack Ukraine Shore Network Malware APT Groups Political Objectives Financial Loss Confidentiality
@@
2018 Port of Barcelona Attack Spain Shore IT Systems Ransomware Unknown N/A Cargo Delay Availability
@@
2021 Death Kitty Ransomware South Africa Shore Network Ransomware Cybercriminal Groups Financial Gain Operations Halted Availability
2021 Attack on Greek Companies Greece Shore Comm System Ransomware Cybercriminal Groups Financial Gain Data Loss Availability
@@
As shown in Table II, maritime cyber incidents have affected numerous countries and regions worldwide, including the United States, Europe (e.g., UK, Germany), Asian-Pacific (e.g., Singapore, Japan, Australia), and Middle East (e.g., UAE). This global distribution of attacks underscores that maritime cybersecurity is an international concern, with no country or region immune to these threats. The cyber attacks documented span a range of categories, including abuse and theft of data, DoS attacks, GPS spoofing and jamming, malware infections (particularly ransomware), social engineering and phishing attempts, targeted attacks, and exploitation of system vulnerabilities. This attack range shows how diverse the cyberthreats that the maritime industry faces are. These attacks are carried out by a similarly varied array of players, including nation-states, hacktivists, cybercriminal organizations, and lone hackers. The threat actors’ identities remained a mystery in numerous instances. This diversity of perpetrators suggests that maritime cyberattacks are motivated by a variety of factors.
III-B Cyber Attack Demonstrations (Academic)
@@
Academic cyber attack demonstrations typically take place in simulated environments or self-developed software that replicate maritime systems and networks. There is a portion of academic research focuses on theoretically feasible attacks by modeling maritime systems to identify potential attack vectors, even if they have not been practically implemented in a full-scale simulation. These demonstrations often focus on specific components of maritime systems
@@
AIS Threats Overview Balduzzi et al. [85] conducted a comprehensive security evaluation of the AIS and categorized these threats into three main types: i) spoofing, ii) hijacking, and iii) availability disruption attacks. Kessler et al. [86] presented a comprehensive analysis of the potential cyber threats to the AIS by drawing upon previous research [87][88][89] that examined the vulnerabilities of Automatic Dependent Surveillance-Broadcast (ADS-B), a similar system used in the aviation industry to provide situational awareness for aircraft. Amro and Gkioulos [90] discussed a novel cyber attack that utilizes the AIS as a covert channel for sending command and control messages and delivering malware to maritime systems.
@@
VDR Threats Overview Harish et al. [91] launched a multi-pronged attack on a commercially available VDR, employing a USB Rubber Ducky for physical infiltration, the Metasploit framework for system exploit, a ransomware simulator to demonstrate data encryption risks, and Nmap to identify exploitable network vulnerabilities. Hopcraft et al. [92] highlighted cybersecurity deficiencies in current VDR systems and standards, and presented they are susceptible to ransomware attacks, malware infections, denial-of-service (DoS) attacks, insider threats and data breaches. Söner et al. [93] used failure modes and effects analysis (FMEA) to assess cyber vulnerabilities and potential attacks against VDR. Data acquisition unit (DAU) is found to be the most vulnerable VDR component, followed by remote access solutions and the bridge control panel. False data injection, command injection, and viruses are ranked as the top three threats in the FMEA risk assessment.
@@
RADAR Threats Overview Junior et al. [94] demonstrated how the attacker could send the command to the ship’s radar via an electronic attack (EA) that generates false radar echoes in a specific pattern. Wolsing et al. [95] presented a taxonomy of possible network cyber attacks on marine RADAR and a simulation environment to implement these attacks. The taxonomy covers denial of service attacks, attacks involving basic image transformations like scaling, rotation and translation, and targeted object manipulations like addition, removal and relocation of radar echoes. Longo et al. [96] demonstrated that it is feasible to develop malware that can autonomously and stealthily target radar systems on ships by exploiting vulnerabilities in standard protocols like NMEA and ASTERIX that lack authentication and encryption. Longo et al. [97] discussed a novel threat model that combines cyber attacks with cyber false flags to target maritime RADAR systems. The proposed cyber attacks manipulate RADAR data to simulate the effects of electronic countermeasures (ECM) like barrage jamming, spot jamming, digital radio frequency memory, and blip enhancement. By making the cyber attacks appear as ECM originating from another ship, it creates a false flag that misleads attribution of the attack. These false flag attacks aim to disrupt radar operations, divert blame, hide the cyber nature of the attack, and project the attacker’s offensive capabilities.
@@
VSAT Threats Overview Santamarta [98] discussed security vulnerabilities found in various widely deployed satellite communications (SATCOM) terminals and described several real-world attack scenarios showing how these flaws could be exploited to leak sensitive military data, interfere with distress communications, spoof navigation data, disable safety systems, and more. Based on the two scenarios presented, VSAT systems in the maritime context face threats of data manipulation, spoofing, and communication disruption. Pavur et al. [99] revealed serious security flaws in how maritime VSAT networks are currently implemented. The insecure VSAT links could allow attackers to modify important data like ECDIS navigational charts and AIS vessel tracking information, impacting ship safety. Under certain conditions, an attacker could hijack active TCP sessions and modify data in real-time.
@@
GNSS Threats Overview Grant et al. [100] conducted practical experiments to validate the threat of GPS jamming on maritime navigation. Bhatti and Humphreys [101] analyzed and demonstrated the ability of an attacker to control a maritime surface vessel by broadcasting counterfeit civil GPS signals. Medina et al. [102] carried out a dedicated jamming measurement campaign on the Baltic Sea in cooperation with the German Federal Network Agency.
@@
ECDIS Threats Overview All five papers written by Svilicic et al. [103, 104, 105, 106, 107] focused on the cyber security aspects of the ECDIS and used similar vulnerability scanning methods. However, each paper had a unique focus. Svilicic et al. [103] presented an estimation of the cyber security vulnerabilities of ECDIS that arise from weaknesses related to the underlying operating system. Svilicic et al. [104] introduced a framework for assessing cyber risks of the ship’s critical systems and assets. Svilicic et al. [105] identified the main threats to the ECDIS system through a survey of the ship’s crew and computational vulnerability scanning using Nessus Professional software. Svilicic et al. [106] highlighted that interconnecting multiple ECDIS workstations with the same vulnerabilities provides an ideal environment for malware to spread. Svilicic et al. [107] emphasized on the cyber security weaknesses of a ECDIS arising from the software’s third-party components. Dyryavyy [108] discussed the cyber security risks and weaknesses within ECDIS and the research revealed that a malicious actor could potentially browse, download, modify, or erase any file on the computer running the ECDIS software [109].
@@
GMDSS Threats Overview While specific GMDSS cyber attack examples are limited, the concepts of jamming, spoofing and DoS [110][111] discussed could all theoretically be used to disrupt GMDSS functionality as ships become more networked and vulnerable to cyber threats. The authors [111] also discussed the risk of malware infections in ship systems. While not specifically mentioning GMDSS, such an attack could potentially affect GMDSS equipment if it’s connected to infected systems.
IV Security Solutions
IV-A Integrated or Holistic
Rather than focusing on individual components, integrated solutions aim to create a comprehensive framework that protects and optimizes the vessel as a whole. For example, an integrated solution might include a centralized security operations center that monitors and manages threats across all vessel systems, from navigation and communication to cargo management and environmental controls. Table III highlights the main holistic security solutions now in use, with an emphasis on contrasting their evaluation techniques and validation datasets.
IV-A1 Machine/Deep Learning-Based Defence
The demonstrated performance and adaptability of some advanced deep learning techniques suggest significant potential for their practical application in safeguarding critical maritime infrastructure against the ever-evolving landscape of cyber threats. In [112], it was revealed that the DLTIF framework exhibited superior performance in automatically extracting and identifying cyber threat patterns compared to traditional machine learning approaches across multiple evaluation metrics. Dual Stack Machine Learning (S2ML) framework is proposed in [113] which takes a novel approach by leveraging entropy-based features extracted from network traffic, a method that proves particularly effective in identifying anomalies indicative of DDoS attacks. Similarly, the authors of [114] leveraged the capabilities of an adaptive incremental passive-aggressive machine learning (AI-PAML) which is designed to continuously update its learning model as new network attacks are detected.
While DLTIF, S2ML and AI-PAML focus on identifying or detecting threats, the Multi-Agent Reinforcement Learning (MARL) [115] approach takes the next step by exploring how AI can be used to actively defend against these threats. The MARL solution establishes a simulation environment for training autonomous cyber defence agents for maritime operational technology (OT) to enhance the resilience of maritime control systems.
IV-A2 Cryptographic Schemes
The distributed nature of maritime systems, with ships, ports, and other infrastructure being geographically dispersed, renders centralized security solutions impractical. Recent research has proposed several cryptographic schemes that provide security in a distributed manner, employing techniques such as symmetric encryption, attribute-based encryption, erasure coding, identity-based encryption, and digital signatures to enable each entity to safeguard its own data and communications. A lightweight authentication protocol [116] using symmetric cryptography primitives like XOR and hash operations can provide mutual authentication and shared key establishment while ensuring vessel privacy. Similarly, an attribute-based secure data aggregation scheme [117] employs constant attributes of maritime terminals for authentication and onboard sensors for data encryption, using zero-knowledge proofs for member certification. Addressing a critical aspect of maritime data security overlooked by the previous schemes, it was proposed that a flexible integrity checking and recovery mechanism [118] is particularly valuable in maritime environments where harsh conditions frequently lead to data corruption. Other prevention methods include an identity-based authentication mechanism [119] that integrates both cryptographic techniques and distributed ledger technologies to ensure not only the confidentiality and integrity of data but also enhance transparency and non-repudiation in maritime communications.
The prevalence of legacy equipment in maritime communication channels, often predating modern security protocols and persisting due to extended lifecycles, presents significant cybersecurity vulnerabilities in today’s interconnected maritime environment. In order to offer a pragmatic strategy for a gradual, cost-effective transition towards more secure maritime communication paradigms, the authors of [120] proposed the SIGMAR framework, which extends the IEC 61162-450 protocol and integrates the Elliptic Curve Digital Signature Algorithm (ECDSA) to provide robust authentication and integrity verification with minimal latency and communication overhead.
IV-A3 Blockchain-Based Defence
A secure communication framework can be harnessed by introducing a private blockchain network integrated with a terrestrial fusion center for authentication purposes [121]. The approach is designed to provide a resilient and trustworthy communication infrastructure for maritime systems, with particular emphasis on unmanned aerial vehicle (UAV)-assisted sensing applications in maritime environments. In [122], a blockchain-based authentication mechanism is utilized to store and validate the identities of vessels and IoT devices. The solution implements a proof-of-authority consensus mechanism and incorporates a threshold-based approach for detecting malicious nodes. Further, blockchain can be used to secure positioning data in maritime environments. A blockchain-based privacy preserving position data sharing approach is leveraged in [123] to address the critical privacy challenges in maritime IoT systems.